# Unlocking Doping Effects on Altermagnetism in MnTe: Emergence of Quasi-altermagnetism

### v2

@@ I Introduction
Our results reveal that, although single-atom substitution lowers the system’s symmetry, it does not destroy altermagnetism-regardless of dopant type or position-due to the preserved six-fold roto-inversion ($S_{6z}$) and mirror ($M_{z}$) symmetry, which maintains the $g$-wave altermagnetism. Pair doping leads to a plethora of possible configurations. Within the constraint of $2\times 2\times 2$ supercell considered here, there exist 120 possible configurations, which can be classified into five symmetry-based families. Among these, three families (46.66%) retain ideal altermagnetism, while the remaining two lack the necessary rotational or roto-inversion mirror symmetries connecting opposite spin sublattices and instead exhibit quasi-altermagnetic behaviour. These quasi-altermagnetic states still show momentum-dependent spin splitting and nearly compensated magnetism. We identify the symmetry-breaking perturbations responsible for the transition from ideal to quasi-altermagnetic states using model Hamiltonian studies. Furthermore, pair-doped structures with lower symmetry exhibit anomalous Hall conductivity (AHC) when magnetization is oriented out-of-plane-a feature absent in pristine MnTe-while a quasi-altermagnetic configuration displays AHC along two different magnetization directions.
@@ II Structural and Computational Details
[…]
@@ III.1 Altermagnetic band structure of pristine MnTe
[…]
@@ III.2 Magnetic structure of MnTe with single non-magnetic atom substitution
Figure 3: Altermagnetic and quasi-altermagnetic states in MnTe for different configurations with a pair of Se substitutions. Substituting a pair of Te atoms with Se atoms in a $\times2$$\times$2 MnTe supercell yields a system with 12.5% doping. This specific substitution results in 120 unique configurations, distributed among five distinct MSGs. The top panel [(a) and (b)] illustrates representative configurations from each MSG, categorized into two classes:(a) those exhibiting ideal altermagnetism, and (b) those displaying quasi-altermagnetic characteristics. The corresponding electronic band structures for these two classes are presented in panels (c) and (d), respectively. The valence and conduction band edges are shown in the insets, where the characteristic features of both altermagnetic and quasi-altermagnetic states are highlighted. The constant energy surfaces for a pair of spin-up (blue) and spin-down channels (red) at $E=E_{F}-0.7$ eV for altermagnets and quasi-altermagnets are shown in panels (e) and (f), respectively. Constant energy surfaces of altermagnets illustrate the six-fold roto-inversion relation of opposite channels, whereas a distortion is visible in that of quasi-altermagnets. Panel (g) provides a pie chart summarizing the distribution of the two-atom-doped configurations across the five MSGs. The grey, green, and yellow segments in the pie chart represent configurations that exhibit ideal altermagnetism, collectively accounting for approximately 46.66% of the total configurations.
@@
Figure 4: Electronic band structures for MnTe supercell with Sb and I pair substitution. Panels (a) and (b) show the band structures of the MnTe supercell in which a pair of non-magnetic atoms have been substituted with Sb, corresponding to different MSGs. Panels (c) and (d) present the band structures for the MnTe supercell with the same substitution by I. The band structures in (a) and (c) exhibit a perfect altermagnetic character, whereas this feature is absent in (b) and (d). Note the shift of the Fermi level due to non-isovalent doping.
@@
Figure 6: Constant energy surfaces of altermagnet and quasi-altermagnet. Panels (a) and (b) show the constant energy surfaces of spin-up (blue) and spin-down (red) bands for pristine MnTe (MSG: $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268)) and Se-doped MnTe (MSG: $C2^{\prime}/m^{\prime}$ (No.12.62)), which is a quasi-altermagnet. The six-fold rotational symmetry connection between the opposite spin bands is clear for altermagnets, while its breaking is evident in the case of Se-doped MnTe, and these do not follow any well-defined symmetry relations.
Figure 7: Anomalous Hall conductivity and Berry curvature of pair doped MnTe with space group $Pm^{\prime}m^{\prime}2$ (No.25.60) and $C2^{\prime}/m^{\prime}2$ (No.12.62). AHC as a function of the Fermi level for Sb-doped MnTe with MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60) and $C2^{\prime}/m^{\prime}2$ (No. 12.62) in panels (a) and (b), and for I-doped cases in panels (c) and (d). Panels (e)–(h) show top views of Berry curvature components $\Omega_{x}$, $\Omega_{y}$, and $\Omega_{z}$ in the Brillouin zone. For Sb, berry curvature is plotted at $E=E_{F}-0.5$ eV, and for I it is at $E=E_{F}+0.15$ eV. For MSG 25.60, only $\sigma_{xy}$ is finite, while other components vanish; for MSG 12.62, $\sigma_{xy}$ and $\sigma_{yz}$ are non-zero and $\sigma_{xz}$ vanishes. Dopant type tunes the AHC magnitude at the Fermi level, while dopant pair positions control the active AHC tensor components.
In Fig. 2(c), we show the band structure of MnTe0.9375Se0.6250, along the altermagnetic $k$-path $\Gamma-LL^{\prime}$, where we observe spin-dependent splitting of the bands in the momentum space, characteristic of altermagnetism. For the visualisation of the altermagnetic behaviour in the full Brillouin zone, in Fig. 2(b), we have plotted the constant energy surface for the spin-up (blue) and spin-down (red) bands at $E=E_{F}-0.85$ eV. The coexistence of compensated magnetization and momentum-dependent spin-splitting confirms that MnTe retains its altermagnetic character even after substituting one Te atom with Se. To investigate whether the specific site of the Te substitution plays any role in determining this behaviour, we performed additional calculations by substituting Se at different Te sites within the supercell. In all cases, the altermagnetic characteristics were preserved, indicating that the observed spin-splitting is intrinsic and independent of the substitution site.
@@
Although magnetic ordering in materials is conventionally studied using magnetic space groups (MSGs), spin space groups (SSGs), where spin and spatial operations are decoupled, have proved to be suitable to characterize altermagnetic materials which exhibit non-relativistic spin-splitting [1, 2, 44]. In the case of MnTe, a collinear altermagnet, spin operations are restricted to either the identity ($\mathcal{E}$) or time-reversal ($C_{2}$), and the symmetry operations connecting opposite spin sublattices are the same in both the SSG and MSG descriptions. Furthermore, since AHC can be explained exclusively within the MSG framework, we employ MSGs for the symmetry analysis in this work.
The space group corresponding to the pristine MnTe is $P6_{3}/mmc$ (No.194), and corresponding MSG is $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268). Maintaining the same supercell geometry, the substitution of one Te atom in MnTe by another element reduces its symmetry, lowering the space group to $P\bar{6}m2$ (No.187). The corresponding MSG is $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). As a result of doping, the lowering of symmetry generates a larger number of symmetry-inequivalent Mn atoms. Since Mn atoms exhibit varying LSM, the spin sublattice structure in the doped system is more complex, as shown in Fig. 2(f). Unlike the pristine case, where the up and down spin sublattices consist of only one Mn atom each, the doped system exhibits two spin sublattices, denoted spin-up in blue and spin-down in red, each containing 8 Mn atoms. These atoms are distributed across two different $c$-planes, with four atoms per plane. The spin-up and spin-down sublattices are related by a mirror symmetry operation ($M_{z}$) passing through the central plane, as illustrated in Fig. 2(f). The mirrored sites have equal and opposite LSM. Furthermore, the elements of spin-up and spin-down sublattices are connected through six-fold rotoinversion: a six-fold rotation about the $z$-axis followed by inversion through the dopant site. To give an example, in Fig. 2(f) and  2(g), we have illustrated how the rotoinversion takes Mn1 to Mn7, Mn2 to Mn5, Mn3 to Mn6, Mn4 to Mn8, and so on. We have further validated that the connecting sites have equal and opposite LSM, leading to a net zero magnetization. Therefore, from the symmetry perspective, this doped system is bound to exhibit altermagnetism. The formation of four nodal planes, three diagonals by rotoinversion ($S_{6z}$) and one basal (M001) by mirror symmetry leads to the emergence of $g$-wave altermagnetism in the doped case.
@@
Next, we examine the electronic and magnetic structure of the doped system with a pair of dopants to gain further insights. While a single dopant results in one configuration irrespective of the position of the dopant in the supercell, a pair of dopants builds a large number of configurations defined by their relative positions. Moreover, a pair of dopants allows us to tailor the symmetry of both geometric and spin space groups, and therefore, we have a wider canvas to further explore the relationship between symmetry and altermagnetism. In the case of MnTe, a pair of dopants at non-magnetic sites gives us 120 possible configurations belonging to five MSG families $C2^{\prime}/m^{\prime}$ (No.12.62), $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), $P\bar{3}m^{\prime}1$ (No.164.89), and $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). In Fig. 3, one representative configurations from each of these five families and the distributions of the five families are shown. A further detailed symmetry analysis shows that, the configurations belonging to $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) have the ideal six-fold roto-inversion, making them eligible to be ideal altermagnets, which is also directly validated from our calculated band structures shown in Fig. 3(c). A closer look at the structure reveals that this configuration has the same symmetry as of the single dopant case. However, we find two more families, namely $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), where the band structures reflect all the characteristic of altermagnetism [see Fig. 3(a)]. Examining these structures further, we find that each of them can be represented through a smaller orthorhombic primitive unit cell with a single dopant in it. These unit cells exhibit two mirror symmetries connecting the opposite spin sublattices, which give rise to $d$-wave altermagnetism with two nodal planes in reciprocal space. It is possible to make a transition from the $d$-wave altermagnetism of the primitive unit cell to the $g$-wave altermagnetism of the supercell, considering that the latter is a hexagonal structure which carries three partially overlapped orthorhombic primitive unit cells (see Appendix A). The two nodal planes from each of the three orthorhombic substructures lead to four distinct nodal planes, which result in $g$-wave altermagnetism.
@@
In the remaining two cases, corresponding to $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89), the situation differs from the three cases discussed earlier. Here, the relative positions of the dopants break the symmetry connecting opposite spin sublattices, as they are no longer related by rotation, translation, or inversion. Therefore, in an ideal context, they are no longer altermagnets. However, band structure analysis for the Se-doped case reveals an interesting observation [Fig. 3(d)]: the band structures exhibit a close resemblance to those of altermagnets, including spin-dependent splitting along the altermagnetic $k$-path, and even though the spin-splitting changes sign, it is not equal on either side of the nodal plane corresponding to the pristine system. These deviations are evident when comparing zoomed views of the valence band maximum and conduction band minimum near $\Gamma$ across all cases. We term these quasi-altermagnetic materials, augmenting the set of ideal altermagnets. Below in a separate subsection, we have comprehensively defined quasi-altermagnetism and explained how it is distinct from ferromagnetism, antiferromagnetism, and altermagnetism (see subsection III.4).
@@
Similar to the band structure, the variations are also evident in the constant energy surface plots of quasi-altermagnets [Fig 3 (f)] compared to that of altermagnets [Fig 3 (e)]. Although the overall momentum-space spin texture of quasi-altermagnets exhibits a qualitative resemblance to the $g$-wave altermagnetic phase, notable differences appears. In the quasi-altermagnetic phase, the constant energy surface exhibits a slight imbalance, where either spin-up or spin-down bands dominate. Moreover, the perfectly matched pairs of spin-up and spin-down bands in altermagnets are absent in quasi-altermagnets.
@@
So far, we have discussed the effect of symmetry lowering induced by non-magnetic atom substitutional doping on altermagnetic properties of MnTe. Substitutional doping at the magnetic atom site also leads to the symmetry lowering, and the resulting configurations exhibit altermagnetism or quasi-altermagnetism depending on the symmetry of the system determined by the relative position of dopant. A discussion regarding this is given in the Appendix  C. This primarily infers that if the roto-inversion is maintained with the doping at the magnetic site, the spin polarized defect bands also show altermagnetic characteristics.

### v3

@@ I Introduction
Our results reveal that, although single-atom substitution lowers the system’s symmetry, it does not destroy altermagnetism-regardless of dopant type or position-due to the preserved six-fold roto-inversion ($S_{6z}$) and mirror ($M_{z}$) symmetry, which maintains the $g$-wave altermagnetism. Pair doping leads to a plethora of possible configurations. Within the constraint of $2\times 2\times 2$ supercell considered here, there exist 120 possible configurations, which can be classified into five symmetry-based families. Among these, three families (46.66%) retain ideal altermagnetism, while the remaining two lack the necessary rotational or roto-inversion mirror symmetries connecting opposite spin sublattices and instead exhibit quasi-altermagnetic behaviour. These quasi-altermagnetic states still show momentum-dependent spin splitting and nearly compensated magnetism. We identify the symmetry-breaking perturbations responsible for the transition from ideal to quasi-altermagnetic states using model Hamiltonian studies. Furthermore, pair-doped structures with lower symmetry exhibit anomalous Hall conductivity (AHC) when magnetization is oriented out-of-plane, a feature absent in pristine MnTe-while a quasi-altermagnetic configuration displays AHC along two different magnetization directions.
@@ II Structural and Computational Details
Figure 2: Altermagnetism in MnTe supercell with single non-magnetic atom substitution. The spin-polarized DOS of the 2 $\times$ 2 $\times$ 2 supercell of MnTe with one Te atom substituted by Se is shown in panel (a), where the blue and red colors represent the spin-up and spin-down channels, respectively. The DOS confirms a magnetically compensated state. Panel (b) shows the constant energy surfaces for the spin-up and spin-down channels (blue and red, respectively) at $E=E_{F}-0.85$ eV, illustrating the spin-splitting in momentum space. Panels (c)–(e) display the spin-resolved electronic band structures of MnTe supercells with a single Te atom substituted by Se, Sb, and I, respectively. All these structures exhibit momentum-dependent spin-splitting along the $\Gamma-LL^{\prime}$ direction. Panels (f) and (g) illustrate symmetry operations connecting opposite spin sublattices in a single Te-substituted MnTe supercell. The mirror plane perpendicular to the $z$-axis (Mz), which connects opposite spin sublattices, is present in panel (f). The structures shown in panels (f) and (g) are related by a six-fold roto-inversion symmetry operation ($S_{6z}$), which consists of a 60° rotation followed by inversion, with the dopant atom serving as the inversion center. Atoms from each sublattice are labelled in planes adjacent to the dopant atom in (f), and their new positions after the $S_{6z}$ operation are shown in panel (g). These features from DFT calculations and symmetry analysis indicate the persistence of $g$-wave altermagnetism in MnTe upon substitution of a single Te atom.
@@ III.1 Altermagnetic band structure of pristine MnTe
Hexagonal MnTe is a well-known $g$-wave altermagnet, characterized by four nodal planes, three parallel to the $z$-axis and one perpendicular to it, as shown in Fig. 1(b). The electronic bands are degenerate along $k$-paths that lie within these nodal planes and become non-degenerate along $k$-paths that lie outside the nodal planes. These non-degenerate bands exhibit AMSS, which varies with momentum and changes sign at opposite momentum directions with respect to a nodal plane. This behavior is illustrated in Fig. 1(c), where the band structure along the $k$-path $\Gamma$-$M$-$K$-$\Gamma$-AA-$L$ (within the nodal plane) is spin-degenerate, while AMSS is observed along $\Gamma-LL^{\prime}$ (outside the nodal planes). Notably, MnTe exhibits a large spin splitting, reaching nearly 1 eV near the valence band maximum. As far as the magnetization is concerned, the Mn atoms of the up (down) sublattice have local spin moments (LSM) of 4.42 $\mu_{B}$ (-4.42 $\mu_{B})$ resulting in zero net magnetization.
@@ III.2 Magnetic structure of MnTe with single non-magnetic atom substitution
Figure 3: Altermagnetic and quasi-altermagnetic states in MnTe for different configurations with a pair of Se substitutions. Substituting a pair of Te atoms with Se atoms in a $\times2$$\times$2 MnTe supercell yields a system with 12.5% doping. This specific substitution results in 120 unique configurations, distributed among five distinct SGs. The top panel [(a) and (b)] illustrates representative configurations from each SGs, categorized into two classes:(a) those exhibiting ideal altermagnetism, and (b) those displaying quasi-altermagnetic characteristics. The corresponding electronic band structures for these two classes are presented in panels (c) and (d), respectively. The valence and conduction band edges are shown in the insets, where the characteristic features of both altermagnetic and quasi-altermagnetic states are highlighted. The constant energy surfaces for a pair of spin-up (blue) and spin-down channels (red) at $E=E_{F}-0.7$ eV for altermagnets and quasi-altermagnets are shown in panels (e) and (f), respectively. Constant energy surfaces of altermagnets illustrate the six-fold roto-inversion relation of opposite channels, whereas a distortion is visible in that of quasi-altermagnets. Panel (g) provides a pie chart summarizing the distribution of the two-atom-doped configurations across the five SGs. The gray, green, and yellow segments in the pie chart represent configurations that exhibit ideal altermagnetism, collectively accounting for approximately 46.66% of the total configurations.
@@
Figure 4: Electronic band structures for MnTe supercell with Sb and I pair substitution. Panels (a) and (b) show the band structures of the MnTe supercell in which a pair of non-magnetic atoms have been substituted with Sb, corresponding to different SGs. Panels (c) and (d) present the band structures for the MnTe supercell with the same substitution by I. The band structures in (a) and (c) exhibit a perfect altermagnetic character, whereas this feature is absent in (b) and (d). Note the shift of the Fermi level due to non-isovalent doping.
@@
Figure 6: Constant energy surfaces of altermagnet and quasi-altermagnet. Panels (a) and (b) show the constant energy surfaces of spin-up (blue) and spin-down (red) bands for pristine MnTe (SG $P6_{3}/mmc$ (No. 194)) and Se-doped MnTe (SG $C2/m$ (No. 12)), which is a quasi-altermagnet. The six-fold rotational symmetry connection between the opposite spin bands is clear for altermagnets, while its breaking is evident in the case of Se-doped MnTe, and these do not follow any well-defined symmetry relations.
Figure 7: Anomalous Hall conductivity and Berry curvature of pair doped MnTe with MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60) and $C2^{\prime}/m^{\prime}$ (No. 12.62). AHC as a function of the Fermi level for Sb-doped MnTe with MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60) and $C2^{\prime}/m^{\prime}$ (No. 12.62) in panels (a) and (b), and for I-doped cases in panels (c) and (d). Panels (e)–(h) show top views of Berry curvature components $\Omega_{x}$, $\Omega_{y}$, and $\Omega_{z}$ in the Brillouin zone. For Sb, berry curvature is plotted at $E=E_{F}-0.5$ eV, and for I it is at $E=E_{F}+0.15$ eV. For MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60), only $\sigma_{xy}$ is finite, while other components vanish; for MSG $C2^{\prime}/m^{\prime}$ (No. 12.62), $\sigma_{xy}$ and $\sigma_{yz}$ are non-zero and $\sigma_{xz}$ vanishes. Dopant type tunes the AHC magnitude at the Fermi level, while dopant pair positions control the active AHC tensor components.
In Fig. 2(c), we show the band structure of MnTe0.9375Se0.6250, along the altermagnetic $k$-path $\Gamma-LL^{\prime}$, where we observe spin-dependent splitting of the bands in the momentum space, characteristic of altermagnetism. For the visualization of the altermagnetic behavior in the full Brillouin zone, in Fig. 2(b), we have plotted the constant energy surface for the spin-up (blue) and spin-down (red) bands at $E=E_{F}-0.85$ eV. The coexistence of compensated magnetization and momentum-dependent spin-splitting confirms that MnTe retains its altermagnetic character even after substituting one Te atom with Se. To investigate whether the specific site of the Te substitution plays any role in determining this behavior, we performed additional calculations by substituting Se at different Te sites within the supercell. In all cases, the altermagnetic characteristics were preserved, indicating that the observed spin-splitting is intrinsic and independent of the substitution site.
@@
Although magnetic ordering in materials is conventionally studied using magnetic space groups (MSGs), spin space groups (SSGs), where spin and spatial operations are decoupled, have proved to be suitable to characterize altermagnetic materials which exhibit non-relativistic spin-splitting [46, 47, 6]. In this work, we analyze the symmetry operations present in the crystallographic space group (SG) of pristine and doped MnTe systems, identifying those that connect opposite spin sublattices and employing SSG [22] to investigate the altermagnetic characteristics from a symmetry perspective. In the later part of the study, MSG [51] is used to discuss the relativistic phenomena, AHE.
The SG corresponding to the pristine MnTe is $P6_{3}/mmc$ (No. 194). Maintaining the same supercell geometry, the substitution of one Te atom in MnTe by another element reduces its symmetry, lowering the SG to $P\bar{6}m2$ (No. 187). As a result of doping, the lowering of symmetry generates a larger number of symmetry-inequivalent Mn atoms. Since Mn atoms exhibit varying LSM, the spin sublattice structure in the doped system is more complex, as shown in Fig. 2(f). Unlike the pristine case, where the up and down spin sublattices consist of only one Mn atom each, the doped system exhibits two spin sublattices, denoted spin-up in blue and spin-down in red, each containing 8 Mn atoms. These atoms are distributed across two different $c$-planes, with four atoms per plane. The spin-up and spin-down sublattices are related by a mirror symmetry operation ($M_{z}$) passing through the central plane, as illustrated in Fig. 2(f). The mirrored sites have equal and opposite LSM. Furthermore, the elements of spin-up and spin-down sublattices are connected through six-fold roto-inversion: a six-fold rotation about the $z$-axis followed by inversion through the dopant site. To give an example, in Fig. 2(f) and  2(g), we have illustrated how the rotoinversion takes Mn1 to Mn7, Mn2 to Mn5, Mn3 to Mn6, Mn4 to Mn8, and so on. We have further validated that the connecting sites have equal and opposite LSM, leading to a net zero magnetization. Therefore, from the symmetry perspective, this doped system is bound to exhibit altermagnetism. The formation of four nodal planes, three diagonals by rotoinversion ($S_{6z}$) and one basal (M001) by mirror symmetry leads to the emergence of $g$-wave altermagnetism in the doped case.
@@
Next, we examine the electronic and magnetic structure of the doped system with a pair of dopants to gain further insights. While a single dopant results in one configuration irrespective of the position of the dopant in the supercell, a pair of dopants builds a large number of configurations defined by their relative positions. Moreover, a pair of dopants allows us to tailor the symmetry of both geometric and spin SGs, and therefore, we have a wider canvas to further explore the relationship between symmetry and altermagnetism. In the case of MnTe, a pair of dopants at non-magnetic sites gives us 120 possible configurations belonging to five SG families, $C2/m$ (No. 12), $Pmm2$ (No. 25), $Fmm2$ (No. 42), $P\bar{3}m1$ (No. 164), and $P\bar{6}m2$ (No. 187). In Fig. 3, one representative configurations from each of these five families and the distributions of the five families are shown. A further detailed spin space symmetry analysis shows that, A-type antiferromagnetic configurations belonging to $P\bar{6}m2$ (No. 187) have the ideal six-fold roto-inversion, making them eligible to be ideal altermagnets, which is also directly validated from our calculated band structures shown in Fig. 3(c). A closer look at the structure reveals that this configuration has the same symmetry as of the single dopant case. However, we find two more families, namely $Pmm2$ (No. 25), $Fmm2$ (No. 42), where the band structures reflect all the characteristics of altermagnetism [see Fig. 3(a)]. Examining these structures further, we find that each of them can be represented through a smaller orthorhombic primitive unit cell with a single dopant in it. These unit cells exhibit two mirror symmetries connecting the opposite spin sublattices, which give rise to $d$-wave altermagnetism with two nodal planes in reciprocal space. It is possible to make a transition from the $d$-wave altermagnetism of the primitive unit cell to the $g$-wave altermagnetism of the supercell, considering that the latter is a hexagonal structure which carries three partially overlapped orthorhombic primitive unit cells (see Appendix A). The two nodal planes from each of the three orthorhombic substructures lead to four distinct nodal planes, which result in $g$-wave altermagnetism.
@@
In the remaining two cases, corresponding to $C2/m$ (No. 12) and $P\bar{3}m1$ (No. 164) the situation differs from the three cases discussed earlier. Here, the relative positions of the dopants break the symmetry connecting opposite spin sublattices, as they are no longer related by rotation, translation, or inversion. Therefore, in an ideal context, they are no longer altermagnets. However, band structure analysis for the Se-doped case reveals an interesting observation [Fig. 3(d)]: the band structures exhibit a close resemblance to those of altermagnets, including spin-dependent splitting along the altermagnetic $k$-path, and even though the spin-splitting changes sign, it is not equal on either side of the nodal plane corresponding to the pristine system. These deviations are evident when comparing zoomed views of the valence band maximum and conduction band minimum near $\Gamma$ across all cases. We term these quasi-altermagnetic materials, augmenting the set of ideal altermagnets. Below in a separate subsection, we have comprehensively defined quasi-altermagnetism and explained how it is distinct from ferromagnetism, antiferromagnetism, and altermagnetism (see subsection III.4).
@@
Similar to the band structure, the variations are also evident in the constant energy surface plots of quasi-altermagnets [Fig 3 (f)] compared to that of altermagnets [Fig 3 (e)]. Although the overall momentum-space spin texture of quasi-altermagnets exhibits a qualitative resemblance to the $g$-wave altermagnetic phase, notable differences appear. In the quasi-altermagnetic phase, the constant energy surface exhibits a slight imbalance, where either spin-up or spin-down bands dominate. Moreover, the perfectly matched pairs of spin-up and spin-down bands in altermagnets are absent in quasi-altermagnets.
@@
So far, we have discussed the effect of symmetry lowering induced by non-magnetic atom substitutional doping on altermagnetic properties of MnTe. Substitutional doping at the magnetic atom site also leads to the symmetry lowering, and the resulting configurations exhibit altermagnetism or quasi-altermagnetism depending on the symmetry of the system determined by the relative position of dopant. A discussion regarding this is given in the Appendix  B. This primarily infers that if the roto-inversion is maintained with the doping at the magnetic site, the spin polarized defect bands also show altermagnetic characteristics.